# Full-Duplex Millimeter Wave MIMO Channel Estimation: A Neural Network Approach

### v1

@@ V Conclusion
In this paper, we studied the channel estimation problem for full-duplex mmWave MIMO systems using NNs. We found that simple NNs with no or few hidden layers can achieve comparable NMSE to deep NNs. The conclusion is the same in the presence of 1-bit ADC distortion, and adding more hidden layers to NNs slightly improves the NMSE of channel estimates. The comparison between the NN and MMSE channel estimators indicated a notable difference. In the high-correlated regime, MMSE is the preferable channel estimator compared to LS and NN, as it can leverage the explicit channel covariance matrix. However, in scenarios where spatial correlation is not extremely high, our results demonstrated that NNs outperform MMSE estimation. Furthermore, the NN-based channel estimator performs remarkably better than MMSE when fewer pilot resources are utilized for channel estimation. This performance allows for a reduction in pilot overhead in full-duplex systems. We applied FNNs to approximate RX-TX channel mapping for the separate antenna configurations in full-duplex systems. Simulation results demonstrated that NNs effectively map channels from downlink UEs to the receive arrays to the channel from the transmit arrays to downlink UEs, particularly in scenarios with high correlation and large antenna arrays. Finally, the complexity analysis revealed that NN-based estimation with fewer than two convolutional hidden layers requires fewer FLOPs compared to MMSE estimation.
Acknowledgment
This work has been supported by the project SEMANTIC, funded by the EU’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 861165.

### v2

@@ V Conclusion
In this paper, we studied the channel estimation problem for full-duplex mmWave MIMO systems using NNs. We show that simple NNs with no or few hidden layers can achieve comparable NMSE to deep NNs. This conclusion holds to some extent even in the presence of few-bit ADC distortion, with additional hidden layers only slightly improving the NMSE of channel estimates. Furthermore, simpler NN architectures with fewer hidden layers demonstrate more robust generalization abilities compared to deep NNs. The comparison between the NN and MMSE channel estimators indicates a notable difference. In the high-correlated regime, MMSE is the preferable channel estimator compared to LS and NN, as it can leverage the explicit channel covariance matrix. However, in scenarios where spatial correlation is not extremely high, our results demonstrate that NNs outperform MMSE estimation. Additionally, the NN-based channel estimator performs remarkably better than MMSE when fewer pilot resources are utilized for channel estimation. This performance allows for a reduction in pilot overhead in full-duplex systems. We applied FNNs to approximate RX-TX channel mapping for the separate antenna configurations in full-duplex systems. Simulation results demonstrate that NNs effectively map channels from downlink UEs to the receive arrays to the channel from the transmit arrays to downlink UEs, particularly in scenarios with high correlation and large antenna arrays. Finally, the complexity analysis reveals that NN-based estimation with fewer than two convolutional hidden layers requires fewer FLOPs compared to MMSE estimation.